OpenStudy (anonymous):

PLEASE simplify: x^2-9/x^2+5x+6

OpenStudy (anonymous):

for one of these problems you need to put in brackets around things so we know what goes with what

OpenStudy (anonymous):

denominator turns out (X+3)(x+2) numerator (x-3)(x+3)

OpenStudy (anonymous):

also, if you post it as a reply then you can use the fancy function options

OpenStudy (anonymous):

so it turrns to be (x-3)/(x+2)

OpenStudy (anonymous):

\[x^2-9/x^2+5x+6\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

tell me which prt is cnfusing ?

OpenStudy (anonymous):

Oh, wait... so the simplified version would be what you posted? Thanks :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(x^2-9)/(X^2+5x+9)) =((x-3)(x+3))/(x+3)(x+2) =(x-3)/(x+2) ;;;(x+3)cancel out